OpenStudy (anonymous):

Your class is conducting an experiment to study geometric sequences. The class forms a line. The first person takes a whole pizza, divides it in half, keeps half, and hands the other half to the next person. The next persons in line do the same thing with their pieces. What portion of the whole pizza does the seventh person in line get? A. 1/256 B. 1/64 C. 1/128 D. 1/32

OpenStudy (anonymous):

not very much \[\frac{1}{2}, \frac{1}{4},\frac{1}{8},\frac{1}{16}\frac{1}{32},\frac{1}{64},\frac{1}{128}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

more easily computed at \(\left(\frac{1}{2}\right)^{7}=\frac{1}{2^7}\)

OpenStudy (anonymous):

yw